OpenStudy (anonymous):

HELP WILL MEDAL AND FAN ON A TIME LIMIT 1 MORE ? Which is a correct comparison of the fractions? A. B. C.

OpenStudy (sushi121212):

Where is the question and answer?

OpenStudy (anonymous):

O SORRY I DON'T KNOW WHAT IS WRONG WITH MY COMPUTER ONE SEC

OpenStudy (anonymous):

15/18 and 25/30 A. 15/18<25/30 B. 15/18>25/30 C. 15/18=25/30

OpenStudy (anonymous):

help

OpenStudy (sushi121212):

The answer is C. 15/18=25/30 because you first have to simplify it: 15 ÷ 3 5 ___ = __ 18 ÷ 3 6 25 ÷5 5 ___ = __ 30 ÷5 6 That shows that they are equal!

OpenStudy (anonymous):

O OK thanks ssssssssssssssssooooooooooooooooooooo much

OpenStudy (anonymous):

GTG BYE